# Debiased Machine Learning and Network Cohesion for Doubly-Robust Differential Reward Models in Contextual Bandits

### 11 Dec 2023 v1

@@ 1 Introduction
At each decision point, a learner receives a context, chooses an action, and observes a reward. The goal is to maximize the expected cumulative reward. High-quality bandit algorithms achieve rewards comparable to those of an optimal policy. To achieve near-optimal performance in mobile health, bandit algorithms must account for (1) the time-varying nature of the outcome variable, (2) nonlinear relationships between states and outcomes, (3) the potential for intervention efficacy to change over time (due, for instance, to habituation as in  Psihogios et al. (2019)), and (4) the fact that similar participants tend to respond similarly to interventions (Künzler et al., 2019).
@@
The paper proceeds as follows. Section 2 summarizes related work. Section 3 describes the model and problem statement. Section 4 describes the algorithm along with the resulting theoretical results. Section 5 describes experimental results for simulations and two mobile health studies. Section 6 concludes with a discussion of limitations and future work.
@@ 3 Model and Problem Statement
where $\boldsymbol{x}(s,a)\in\mathbb{R}^{p\times 1}$ is a vector of features of the state and action, $\delta_{a>0}$ is an indicator that takes value $1$ if $a>0$ and 00 otherwise, and $g_{t}(s)$ is a baseline reward that observed when individuals are randomized to not receive any treatment. This can be an arbitrary, potentially nonlinear function of state $s$ and time $t$. Equation (1) is equivalent to assuming a linear differential reward for any $a>0$; i.e., $\Delta_{i,t}(s,a):=r_{i,t}(s,a)-r_{i,t}(s,0)$ is linear in $\boldsymbol{x}(s,a)$, whose parameter $\theta_{i,t}\in\mathbb{R}^{p}$ is allowed to depend both on the individual $i$ and time $t$.
@@ 3.1 DML and Doubly Robust Differential Reward
We refer to these $\Delta_{i,t}(s_{i,t},\bar{a}_{i,t})$ as the differential reward. Below, we define a pseudo-reward with the same expectation in reference to pseudo-outcomes from the causal inference literature (Bang & Robins, 2005; Kennedy, 2020). Let $f_{i,t}(s,a)$ be a working model for the true conditional mean $r_{i,t}(s,a)$. Then, following connections to pseudo-outcomes and doubly-robust (DR) estimators (Kennedy, 2020; Shi & Dempsey, 2023), we define the pseudo-reward $\tilde{R}_{i,t}^{f}(s,\bar{a})$ given state $S_{i,t}=s$ and potential arm $\bar{a}$
@@ 3.2 Nearest Neighbor Regularization
Above, we considered a single individual $i$ under a time-invariant linear differential reward function; i.e., $\boldsymbol{x}(s_{i,t},a_{i,t})^{\top}\theta$ where $\theta\in\mathbb{R}^{p}$. Here, we consider the setting of $N$ independent individuals and a time-invariant linear differential reward with individual-specific parameter; i.e., $\theta_{i}\in\mathbb{R}^{p}$. If $\theta_{i}$ were known a priori, then one could construct a network based on $L_{2}$-distances $\{d(i,j):=\|\theta_{i}-\theta_{j}\|_{2}^{2}\}_{j\not=i}$.
@@
We define the Laplacian via the $N\times MN$ incidence matrix $B$. The element $B_{v,e}$ corresponds to the $v$-th vertex (user) and $e$-th edge. Denote the vertices of $e$ as $v_{i}$ and $v_{j}$ with $i$ $j$. $B_{ve}$ is then equal to 1 if $v=v_{i}$, -1 if $v=v_{j}$, and 0 otherwise. The Laplacian matrix is then defined as $L=BB^{\top}$. We can then adapt (4) by summing over participants and including a network cohesion penalty similar to Yang et al. (2020):
@@
where $\Theta:=(\theta_{1},\ldots,\theta_{N})^{\top}\in\mathbb{R}^{N\times p}$. The penalty is small when $\theta_{i}$ and $\theta_{j}$ are close for connected users. Following Assumption 2 and above discussion, we further assume:
@@ 4.1 Algorithm
By performing a joint asymptotic analysis with respect to the total number of individuals ($N$) and time points ($T$), we can relax the assumption of a single time-invariant linear advantage function and allow $\theta_{i,t}\in\mathbb{R}^{p}$ to depend on both the individual $i$ and time $t$. Here, we let $\theta_{i,t}=\theta+\theta_{i}^{\textrm{ind}}+\theta_{t}^{\textrm{time}}$; i.e., include (i) an individual-specific, time-invariant term $\theta_{i}$, and (ii) a shared, time-specific term $\theta_{t}$. This setup is similar to the intelligentpooling method of Tomkins et al. (2021); however, rather than assume individuals and time points are unrelated iid samples, we assume knowledge of some network information (e.g., the similarity of certain individuals or proximity in time) and regularize these parameters accordingly to ensure network cohesion.
The DML-TS-NNR algorithm is shown in Algorithm 1. To see the motivation, consider the following. We first assume that we have access to two nearest neighbor graphs, $G_{\text{user}}$ and $G_{\text{time}}$, where each characterizes proximity in the user- and time-domains respectively. Then at at stage $k$, we estimate all parameters, e.g. $\Theta_{k}=\textrm{vec}[\left(\theta,\theta_{1}^{\textrm{ind}},\ldots,\theta_{k}^{\textrm{ind}},\theta_{1}^{\textrm{time}},\ldots,\theta_{k}^{\textrm{time}}\right)]\in\mathbb{R}^{p(2k+1)}$ , by minimizing the following penalized loss function $L_{k}(\Theta_{k};\lambda,\gamma)$, which is defined as the following expression:
@@
where $v_{k}^{2}\equiv Cc\log^{2m}(k)k^{-1/2}+\sigma^{2}c^{2}$
@@ 4.2 Regret Analysis
This is a version of pseudo-regret Audibert et al. (2003). Compared to standard regret, the randomness of the pseudo-regret is due to $\{A_{i,t}\}_{t=1}^{\top}$ since the error terms $\{\epsilon_{i,t}\}_{t=1}^{\top}$ are removed in the definition.

### 15 Dec 2023 v2

@@ 1 Introduction
At each decision point, a learner receives a context, chooses an action, and observes a reward. The goal is to maximize the expected cumulative reward. High-quality bandit algorithms achieve rewards comparable to those of an optimal policy. To achieve near-optimal performance in mHealth, bandit algorithms must account for (1) the time-varying nature of the outcome variable, (2) nonlinear relationships between states and outcomes, (3) the potential for intervention efficacy to change over time (due, for instance, to habituation as in  Psihogios et al. (2019)), and (4) the fact that similar participants tend to respond similarly to interventions (Künzler et al., 2019).
@@
The paper proceeds as follows. Section 2 summarizes related work. Section 3 describes the model and problem statement. Section 4 describes the algorithm along with the resulting theoretical results. Section 5 describes experimental results for simulations and two mHealth studies. Section 6 concludes with a discussion of limitations and future work.
@@ 3 Model and Problem Statement
where $\boldsymbol{x}(s,a)\in\mathbb{R}^{p\times 1}$ is a vector of features of the state and action, $\delta_{a>0}$ is an indicator that takes value $1$ if $a>0$ and 00 otherwise, and $g_{t}(s)$ is a baseline reward that is observed when individuals are randomized to not receive any treatment. This can be an arbitrary, potentially nonlinear function of state $s$ and time $t$. Equation (1) is equivalent to assuming a linear differential reward for any $a>0$; i.e., $\Delta_{i,t}(s,a):=r_{i,t}(s,a)-r_{i,t}(s,0)$ is linear in $\boldsymbol{x}(s,a)$, whose parameter $\theta_{i,t}\in\mathbb{R}^{p}$ is allowed to depend both on the individual $i$ and time $t$.
@@ 3.1 DML and Doubly Robust Differential Reward
We refer to this $\Delta_{i,t}(s_{i,t},\bar{a}_{i,t})$ as the differential reward. Below, we define a pseudo-reward with the same expectation in reference to pseudo-outcomes from the causal inference literature (Bang & Robins, 2005; Kennedy, 2020). Let $f_{i,t}(s,a)$ be a working model for the true conditional mean $r_{i,t}(s,a)$. Then, following connections to pseudo-outcomes and doubly-robust (DR) estimators (Kennedy, 2020; Shi & Dempsey, 2023), we define the pseudo-reward $\tilde{R}_{i,t}^{f}(s,\bar{a})$ given state $S_{i,t}=s$ and potential arm $\bar{a}$
@@ 3.2 Nearest Neighbor Regularization
Above, we considered a single individual $i$ under a time-invariant linear differential reward function; i.e., $\boldsymbol{x}(s_{i,t},a_{i,t})^{\top}\theta$ where $\theta\in\mathbb{R}^{p}$. Here, we consider the setting of $N$ independent individuals and a time-invariant linear differential reward with an individual-specific parameter; i.e., $\theta_{i}\in\mathbb{R}^{p}$. If $\theta_{i}$ were known a priori, then one could construct a network based on $L_{2}$-distances $\{d(i,j):=\|\theta_{i}-\theta_{j}\|_{2}^{2}\}_{j\not=i}$.
@@
We define the Laplacian via the $N\times MN$ incidence matrix $B$. The element $B_{v,e}$ corresponds to the $v$-th vertex (user) and $e$-th edge. Denote the vertices of $e$ as $v_{i}$ and $v_{j}$ with $i>j$. $B_{v,e}$ is then equal to 1 if $v=v_{i}$, -1 if $v=v_{j}$, and 0 otherwise. The Laplacian matrix is then defined as $L=BB^{\top}$. We can then adapt (4) by summing over participants and including a network cohesion penalty similar to Yang et al. (2020):
@@
where $\Theta:=(\theta_{1},\ldots,\theta_{N})^{\top}\in\mathbb{R}^{N\times p}$. The penalty is small when $\theta_{i}$ and $\theta_{j}$ are close for connected users. Following Assumption 2 and the above discussion, we further assume:
@@ 4.1 Algorithm
By performing a joint asymptotic analysis with respect to the total number of individuals ($N$) and time points ($T$), we can relax the assumption of a single time-invariant linear advantage function and allow $\theta_{i,t}\in\mathbb{R}^{p}$ to depend on both the individual $i$ and time $t$. Here, we let $\theta_{i,t}=\theta+\theta_{i}^{\textrm{ind}}+\theta_{t}^{\textrm{time}}$; i.e., we include (i) an individual-specific, time-invariant term $\theta_{i}$, and (ii) a shared, time-specific term $\theta_{t}$. This setup is similar to the intelligentpooling method of Tomkins et al. (2021); however, rather than assume individuals and time points are unrelated i.i.d. samples, we assume knowledge of some network information (e.g., the similarity of certain individuals or proximity in time) and regularize these parameters accordingly to ensure network cohesion.
The DML-TS-NNR algorithm is shown in Algorithm 1. To see the motivation, consider the following. We first assume that we have access to two nearest neighbor graphs, $G_{\text{user}}$ and $G_{\text{time}}$, where each characterizes proximity in the user- and time-domains, respectively. Then at stage $k$, we estimate all parameters, e.g. $\Theta_{k}=\textrm{vec}[\left(\theta,\theta_{1}^{\textrm{ind}},\ldots,\theta_{k}^{\textrm{ind}},\theta_{1}^{\textrm{time}},\ldots,\theta_{k}^{\textrm{time}}\right)]\in\mathbb{R}^{p(2k+1)}$ , by minimizing the following penalized loss function $L_{k}(\Theta_{k};\lambda,\gamma)$, which is defined as the following expression:
@@
where $v_{k}^{2}$ is defined in Corollary 4
@@ 4.2 Regret Analysis
This regret is a version of pseudo-regret Audibert et al. (2003). Compared to standard regret, the randomness of the pseudo-regret is due to $\{A_{i,t}\}_{t=1}^{\top}$ since the error terms $\{\epsilon_{i,t}\}_{t=1}^{\top}$ are removed in the definition.